Suppose the scores on a statistic exam are normally distributed with a mean of 77 and a variance of 25.

  1. What is the 25th percentile of the scores?

  2. What is the percentile of someone who got a score of 62?

  3. What proportion of the scores are between 80 and 90?

  4. Suppose you select 35 tests at random, what is the proportion of scores above 85?
Answer #1

Let X be the score on that statistic exam.Then X~ N(77,25)

1.25th percentile of the population is 25th percentile of N(77,25) which is 73.627

2. percentile for 62 is P(X<=62) *100 = 0.0013 *100 = 0.13th percentile

3. required answer is P(80<=X<=90) = 0.2696

4. the proportion will be same as P(X>85)*100 = 0.0548*100 = 5.48%